Solve the system by using substitution. 2x+2y= -6, 3x-2y= -14
Answer by mananth(16946)

2 x + 2 y = -6 .............1
-2 y = -14 -3 x ---------2
/ -2 =
y = 7 + 1.5 x
Substitute y in (1)
2 x + 2 ( 7 + 1.5 x ) = -6
2 x + 14 + 3 x = -6
2 x + 3 x = -6 -14
5 x = -20
/ 5
x= -4
Plug the value of x in (1)
2 x + 2 y = -6
2 * -4 + 2 y = -6
-8 + 2 y = -6
2 y = 2
/ 2
y= 1
m.ananth@hotmail.ca